OpenStudy (aonz):

Hi guys :) represent the solution set of the inequality on a number line. |x^2-1|<= 4

OpenStudy (aonz):



OpenStudy (anonymous):

\[ |x^2-1|<= 4\implies -4<=x^2-1<= 4 \]

OpenStudy (anonymous):

Which gives us: \[ 0\leq x^2+3 \leq 8 \]And \[ -8\leq x^2-5\leq 0 \]

OpenStudy (anonymous):

Find roots in both cases.